OpenStudy (anonymous):

Find the length of the radius of the following circle. (x + 2)² + y² = 10 5 √(10) 10

OpenStudy (anonymous):

(x - h)^2+(y-k)^2=r^2 is the equation of a circle, to find the radius (r) just take the square root of r \[\sqrt{10}\]

OpenStudy (anonymous):

Standard form of a circle: \[(x-h)^{2}+(y-k)^{2}=r^{2} \] Your circle: \[(x-(-2))^{2}+(y-0)^{2}= (\sqrt{10})^{2}\] So: \[r=\sqrt{10}\]